OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 6 2x - 4y = 10

OpenStudy (anonymous):

See, make either coefficient of \(x\) or \(y\) just equal to what is that in second equation..

OpenStudy (anonymous):

Firstly, multiply first equation by \(2\)..

OpenStudy (blake57roger):

"If you multiply equation one by 2 you get 2x-4y=12 2x-4y=10" Since 2x-4y=2x-4y, that implies that 12=10, which is false, so there is no solution. The lines are parallel.

OpenStudy (anonymous):

there is a solution

OpenStudy (anonymous):

So, you will get two equations as : \[2x - 4y = 12 \\ 2x - 4y = 10\]

OpenStudy (blake57roger):

Method two: x-2y=6 --> y=x/2-3 2x-4y=10 --> y=x/2-5/2 Both lines have the same slope but a different intercept, ergo, parallel, and no solution.

OpenStudy (anonymous):

yes^

OpenStudy (anonymous):

As, at the same time, \((2x-4y)\) cannot be equal to \(10\) and \(12\) both, so this means lines are parallel and it is having NO SOLUTION..

OpenStudy (anonymous):

@blake57roger is right

OpenStudy (blake57roger):

Thanks